The number of employees,average daily wages per employee and varaiance of daily wages per employee for two factory given below
                                Factory a  factory b
No of employee                50        100
Average of employee      120       85
Variance of daily wages  9          16
In which factory is there greter variation in the distribution of daily wages per employee?
 

Grade:12th pass

1 Answers

Bhagyashri verma
5 years ago
S.D for factory A = \sqrt Variance =3 
Coefficient of variation =  (SD/ Avg) * 100 = 3/120 * 100 = 2.5 %
 
S.D for factory A = \sqrt Variance =4
 
Coefficient of variation =  (SD/ Avg) * 100 = 4/85 * 100 = 4.75%
 
Hence greater variation in distribution of wages in factory B